### v1

# Statistical features of quantum chaos using the Krylov operator complexity

Abstract
Recently the Krylov operator complexity is proposed to evaluate the operator growth in quantum systems, and the variance of its Lanzcos coefficients is used as an important parameter for chaos. In this paper, we generate samples of random initial operators from given probability distribtions (GOE, GUE and the uniform distribution). For the Sinai billiard model, we study the statistical properties of the variance of Lanzcos coefficients in the associated Krylov space. Depending on whether the system is chaotic or not, the resulting distribution of the variance have different behaviors. In the nonchaotic case, the resulting distributions are almost overlapping together. In the chaotic case, they split into two well-separated groups. Besides, all the resulting distributions of the variance are the normal distribution, as long as the matrix size of the initial operator is large enough. This overlap-to-separation behavior, in the resulting distribution of the variance of Lanczos coefficients, may be a characteristics to distinguish chaotic and nonchaotic dynamics.
1 Introduction
The physics of complexity is an important topic that contributes to our understanding of both classical and quantum chaos, especially the growth of operator complexity. Recently the Krylov complexity Parker:2018yvk is proposed as an upper bound on the operator complexity, including the famous bound $\lambda_{L}\leq 2\pi k_{B}T/\hbar$ Maldacena:2015waa of the growth of the out-of-time-order correlation function (OTOC) 1969JETP28:1200L . The operator dynamics in the Krylov space resembles the hopping dynamics on the 1d chain Parker:2018yvk . From this, the connection between the dynamics of operator complexity and the Anderson localization is studied in Rabinovici:2021qqt , where the variance $\sigma^{2}$ of the Lanczos coefficients is connected to the late-time saturation of Krylov complexity.
Then a statistical correlation between the Krylov complexity and the level statistics Dyson:1962JMP ; Gutzwiller:1971fy ; Berry:1977RSPSA ; Bohigas:1983er ; Blumel:1990PhRvL are firstly reported in spin systems Rabinovici:2021qqt ; Rabinovici:2022beu and billiard systems Hashimoto:2023swv . For the billiard system Hashimoto:2023swv , the variance of Lanczos coefficents is computed from the Krylov space of the momentum operator, and the level statistics are computed as usual. As the system Hamiltonian varies from nonchaotic to chaotic behavior, a sequence of data are generated for the level statistics and the variance of Lanczos coefficients. The statistical correlation coefficient between these two sequences of data are not zero. For the chaotic spin system Rabinovici:2022beu , the late-time saturation of Krylov complexity approaches the behavior of Random Matrix Theory (RMT) under certain circumstances. Also, the operator generating the Krylov space is randomly selected and three different operators are compared Rabinovici:2022beu . Recently in Craps:2024suj , the Krylov complexity is studied by choosing multiple initial operators, instead of one single operator.
Along this line, it is a natural and straightforward question to ask what would happen if we do statistics on initial operators over a given probability distribution. In this paper, we study this possibility. We generate samples of initial operators from a probability distribution (GOE, GUE and the uniform distribution) and study the resulting distribution of the variance $\sigma^{2}$ of Lanzcos coefficients. Interestingly, we find that for the nonchaotic case, the resulting distributions of the variance $\sigma^{2}$ are nearly the same for different distributions of initial operators. And surprisingly, for the chaotic case, the resulting distributions of the variance $\sigma^{2}$ are obviously different. More intuitively, for different distributions of initial operators, the resulting plots of distributions are overlapped together in the nonchaotic case, and divide into two well-separated groups in the chaotic case.
In detail, we consider a typical billiard system, the Sinai billiard, which originates from a quantization of classical chaos BERRY1981163 . The computation is done in the energy representation and operators become matrices. For the random initial operator $\mathcal{O}$ coming from the probability distribution GOE (GUE and the uniform distribution), the corresponding random matrix is sampled from the standard GOE (GUE and Hermitian matrices whose matrix elements follow the uniform distribution). Then the variance $\sigma^{2}$ of Lanczos coefficients is obtained in the associated Krylov space, when the system is chaotic or nonchaotic. By sampling the initial operator many times for each distribution, we can analyze the resulting statistical distribution of the variance $\sigma^{2}$. We find that the resulting distribution of $\sigma^{2}$ is almost always the normal distribution, independent of the initial operator distribution. As the system changes from nonchaotic to chaotic, the resulting normal distributions of $\sigma^{2}$ is computed and the logarithm of its statistical average $\langle\sigma^{2}\rangle$ and standard deviation are shown in Figure 5. In the nonchaotic case, the resulting distributions of $\sigma^{2}$ are almost overlapping together. In the chaotic case, they split into two groups well-separated from each other: one group consists of GOE, the uniform distribution of pure real numbers (URE) and of pure imaginary numbers (UIM), the other group is of GUE and the uniform distribution of complex numbers (UCP).
The paper is organized as follows. In Section 2, we briefly review the Krylov operator space. In Section 3, we explain the sampling of the random initial operator. In Section 4, we give a qualitative explanation for the reason why the resulting distribution of the variance $\sigma^{2}$ is always the normal distribution. In Section 5, we show the detailed results of the model-the Sinai billiard. In Section 6, we discuss the physics of our results and conclude with a discussion of open questions.
2 Preliminary: Krylov operator space
Here we briefly review the algorithms of Krylov operator space Parker:2018yvk . Readers familiar with it can skip this section. For a quantum system with the Hamiltonian $H$, the time evolution of a given operator $\mathcal{O}$ in the Heisenberg picture is given by
$\mathcal{O}(t)=e^{iHt}\mathcal{O}(0)e^{-iHt}$ (2.1)
Using the Baker-Campbell-Hausdorff formula, it can be expanded as
$\mathcal{O}(t)=\sum_{n=0}^{\infty}\frac{(it)^{n}}{n!}\mathcal{L}^{n}\mathcal{O}(0),$ (2.2)
where :-$\mathcal{L}\coloneq[H,\cdot]$ is the Liouvillian superoperator. This sequence of operators $\mathcal{O}(0)$, $\mathcal{L}\mathcal{O}(0)$, $\mathcal{L}^{2}\mathcal{O}(0)$, $\cdots$ spans a space associated with the operator $\mathcal{O}(t)$. This space is the Krylov subspace generated by $\mathcal{L}$ (and hence $H$) and $\mathcal{O}(0)$, that is,
$\mathcal{K}_{m}(\mathcal{L},\mathcal{O}(0))\equiv\left\{\mathcal{O}(0),\mathcal{L}\mathcal{O}(0),\mathcal{L}^{2}\mathcal{O}(0),\cdots,\mathcal{L}^{m}\mathcal{O}(0)\right\}$ (2.3)
An orthonormal basis, called the Krylov basis, can be constructed for $\mathcal{K}_{m}(\mathcal{L},\mathcal{O}(0))$. This construction also generates a sequence of important nubmers called the Lanczos coefficients, because the algorithm used is the Lanczos algorithm( or similar algorithms). The orthogonality is guaranteed by the inner product $(A|B)=\operatorname{tr}[A^{\dagger}B]$, which can be considered as a simplification of the Wightman inner product at infinite temperature. The Lanczos algorithm can be briefly stated as following:
1:$\|\mathcal{O}\|=\sqrt{(\mathcal{O}|\mathcal{O})}$
2:$b_{0}\leftarrow 0$
3:$\mathcal{O}_{-1}\leftarrow 0$
4:$\mathcal{O}_{0}\leftarrow\mathcal{O}(0)/\|\mathcal{O}(0)\|$
5:$n\leftarrow 1$
6:while $b_{n}\neq 0$ do
7:     $\mathcal{A}_{n}\leftarrow\mathcal{L}\mathcal{O}_{n-1}-b_{n-1}\mathcal{O}_{n-2}$
8:     $b_{n}\leftarrow\|\mathcal{A}_{n}\|$
9:     $\mathcal{O}_{n}\leftarrow\mathcal{A}_{n}/b_{n}$
10:     $n\leftarrow n+1$
11:end while
The Krylov basis is given by $\left\{\left|O_{0}\right),\left|O_{1}\right)\cdots\left|O_{n_{k}}\right)\right\}$ and the dimension of Krylov operator space is $n_{k}=dim(\mathcal{K}_{m})-1$. The sequence of positive numbers $\left\{b_{n}\right\}$ is called the Lanczos coefficient and is essentially the matrix element
$L_{nm}\equiv(\mathcal{O}_{n}|\mathcal{L}|\mathcal{O}_{m})=\begin{pmatrix}0&b_{1}&0&0&\cdots\\ b_{1}&0&b_{2}&0&\cdots\\ 0&b_{2}&0&b_{3}&\cdots\\ 0&0&b_{3}&0&\ddots\\ \vdots&\vdots&\vdots&\ddots&\ddots\end{pmatrix}$ (2.4)
The Lanczos coefficient $\left\{b_{n}\right\}$ contains all the information about the evolution of the operator. To see this, firstly expand the operator $\mathcal{O}(t)$ in terms of the Krylov basis
$\mathcal{O}(t)=\sum_{n=0}^{n_{k}-1}i^{n}\varphi_{n}(t)\mathcal{O}_{n},\quad\varphi_{n}(t)\equiv(\mathcal{O}_{n}|\mathcal{O}(t))/i^{n},$ (2.5)
and then substitute it into the Heisenberg equation. This would give
$\dot{\varphi}_{n}(t)=b_{n}\varphi_{n-1}(t)-b_{n+1}\varphi_{n+1}(t),$ (2.6)
where the dot is the time derivative and the initial condition is $\varphi_{n}(0)=\delta_{n0}\|O\|$. The logarithmic variance $\sigma$ of the Lanczos coefficient $\left\{b_{n}\right\}$ is defined by
$\sigma^{2}\coloneqq\text{Var}(x_{i})=\left\langle x^{2}\right\rangle-\left\langle x\right\rangle^{2},x_{i}=\ln{\frac{b_{2i-1}}{b_{2i}}}.$ (2.7)
In Parker:2018yvk , the equation (2.6) is interpreted as the dynamincs on a 1d chain, with $\varphi_{n}(t)$ being the wave function and $b_{n}$ being hopping amplitude. Then the Krylov space can be associated with systems of spin chains. In this analog, the connection between the dynamics of complexity and the Anderson localization is studied in Rabinovici:2021qqt , and the importance of the variance of Lanczos coefficient is recognized.
3 Random initial Operators
In this paper, we choose samples of random initial operators from different probability distributions and study the resulting distribution of the variance $\sigma^{2}$ of the Lanczos coefficients. Specifically, we consider random initial operators ($\mathcal{O}_{GOE}$, $\mathcal{O}_{GUE}$ and $\mathcal{O}_{uni}$) from the Gaussian orthogonal ensemble (GOE), the Gaussian unitary ensemble (GUE) and the uniform distribution.
These operators are evaluated in the energy representation. In the case of GOE, the random matrix, $\mathcal{O}_{mn}\coloneqq\left\langle m\right|\mathcal{O}_{GOE}\left|n\right\rangle$, is a real symmetric matrix whose matrix elements satisfy the following joint probability density function mehta1991random :
$P\left(\mathcal{O}_{11},\mathcal{O}_{12},\cdots,\mathcal{O}_{NN}\right)=(\frac{1}{2\pi})^{\frac{N}{2}}(\frac{1}{\pi})^{\frac{N^{2}-N}{2}}\exp\left[-\frac{1}{2}\text{Tr}(\mathcal{O}_{GOE}^{2})\right].$ (3.1)
In the case of GUE, the random matrix, $\mathcal{O}^{GUE}_{mn}\coloneqq\left\langle m\right|\mathcal{O}_{GUE}\left|n\right\rangle$ is Hermitian whose elements satisfy the following joint probability density function:
$P\left(\mathcal{O}^{(0)}_{11},\mathcal{O}^{(0)}_{12},\cdots,\mathcal{O}^{(0)}_{NN}\right)=(\frac{1}{2\pi})^{\frac{N}{2}}(\frac{1}{\pi})^{\frac{N^{2}-N}{2}}\exp\left[-\frac{1}{2}\text{Tr}(\mathcal{O}^{(0)2})\right]$ (3.2)
$P\left(\mathcal{O}^{(1)}_{11},\mathcal{O}^{(1)}_{12},\cdots,\mathcal{O}^{(1)}_{NN}\right)=(\frac{1}{2\pi})^{\frac{N}{2}}(\frac{1}{\pi})^{\frac{N^{2}-N}{2}}\exp\left[-\frac{1}{2}\text{Tr}(\mathcal{O}^{(1)2})\right]$ (3.3)
with $\mathcal{O}^{(0)}$ being the real part of $\mathcal{O}^{GUE}_{mn}$, and $\mathcal{O}^{(1)}$ being the imaginary part of $\mathcal{O}^{GUE}_{mn}$. In the case of uniform distribution, three types of random operators are studied. Their random matrices belong to three types: purely real symmetric (URE), purely imaginary Hermitian (UIM), and complex Hermitian (UCP).
For convenience, we introduce the distribution function $f_{\sigma^{2}}(\mathcal{O})$, whose value is the variance $\sigma^{2}$ of the Lanczos coefficients generated by the Krylov subspace $\mathcal{K}_{m}(\mathcal{L},\mathcal{O})$ of the initial operator $\mathcal{O}$
$f_{\sigma^{2}}(\mathcal{O})\coloneqq\text{the variance }\sigma^{2}\text{ of Lanczos coefficients of the Krylov subspace }\mathcal{K}_{m}(\mathcal{L},\mathcal{O}).$ (3.4)
Obviously, the resulting distribution function $f_{\sigma^{2}}$ depends on the initial distribution of the operator $\mathcal{O}$. Interestingly, when compare different initial probability distributions of $\mathcal{O}$, the behavor of the resulting distribution function $f_{\sigma^{2}}$ depends on whether the system is nonchaotic or chaotic: nearly complete overlapping $f_{\sigma^{2}}$ in the nonchaotic case, and two well-separated groups of $f_{\sigma^{2}}$ in the chaotic case. For all cases, the distribution function $f_{\sigma^{2}}$ is a normal distribution.
4 Statistical properties of Lanczos coefficients
Before showing detailed results, we give a qualitative analysis of the reason why the resulting distributions are normal distributions. This is related with the nature of the Krylov algorithm itself and can be understood using the central limit theorem of probability theory, so we explain it here in this section.
Firstly, we prove the following lemma by mathematical induction, which essentially says that the matrices involved in the Krylov algorithm depends only on the energy levels and the initial random matrix.
Lemma 1.
Matrix elements $\mathcal{O}^{(k)}_{mn}$ of Krylov basis $\mathcal{O}_{k}$ in the energy representation can be written as:
$\mathcal{O}^{(k)}_{mn}=C^{(k)}_{mn}(b_{1},b_{2},\cdots,b_{k};E_{mn})O_{mn}^{(0)}$ (4.1)
where $E_{mn}\coloneqq E_{m}-E_{n}$.
Proof.
$k=0,1$ are clearly true, because
$\mathcal{O}^{(0)}_{mn}=\mathcal{O}^{(0)}_{mn}$ (4.2)
$\mathcal{O}^{(1)}_{mn}=\frac{1}{b_{1}}\mathcal{A}_{mn}^{(1)}=\frac{1}{b_{1}}(E_{mn}\mathcal{O}^{(0)}_{mn}),b_{1}=\|E_{mn}\mathcal{O}^{(0)}_{mn}\|$ (4.3)
Assume that $k-1,k-2$ holds, then $k$ also holds, because
$\mathcal{A}_{mn}^{(k)}$ $=E_{mn}\mathcal{O}^{(k-1)}_{mn}-b_{k-1}\mathcal{O}^{(k-2)}_{mn}$ (4.4)
$=[E_{mn}C^{(k-1)}_{mn}(b_{1},b_{2},\cdots,b_{k-1};E_{mn})-b_{n-1}C^{(k-2)}_{mn}(b_{1},b_{2},\cdots,b_{k-2};E_{mn})]O_{mn}^{(0)}$ (4.5)
$=\tilde{C}^{(k)}_{mn}(b_{1},b_{2},\cdots,b_{k-1};E_{mn})O_{mn}^{(0)}$ (4.6)
and
$b_{k}$ $=\|\tilde{C}^{(k)}_{mn}(b_{1},b_{2},\cdots,b_{k-1};E_{mn})O_{mn}^{(0)}\|$ (4.7)
$\mathcal{O}^{(k)}_{mn}$ $=\frac{1}{b_{k}}\mathcal{A}_{mn}^{(k)}=C^{(k)}_{mn}(b_{1},b_{2},\cdots,b_{k};E_{mn})O_{mn}^{(0)}.$ (4.8)
Then by mathematical induction the lemma is proved. ∎
From this lemma, we see that all Lanczos coefficients $b_{k}=b_{k}(O_{mn}^{(0)};E_{mn})$ depend only on the energy levels and the initial random operator $O_{mn}^{(0)}$. In practical computation, the energy levels are truncated to $\mathcal{N}_{max}$ and the random matrix $\mathcal{O}^{(0)}_{mn}$ is of size $\mathcal{N}_{max}\times\mathcal{N}_{max}$. If the value of $\mathcal{N}_{max}$ is large, we would have large-enough random numbers, so qualitatively the resulting distribution can be estimated by the central limit theorem. This is one aspect related to the large value of $\mathcal{N}_{max}$.
In the other aspect, suppose the variance $\sigma_{\mathcal{O}}^{2}$ is computed from $2K$ Lanczos coefficients of a given initial operator $\mathcal{O}^{(0)}$, with $2K$ being the algorithm-steps selected from the Krylov space. By definition (2.7), we have
$\sigma_{\mathcal{O}}^{2}$ $=\frac{1}{K}\sum_{i=1}^{K}\bigg{(}x_{i}(\mathcal{O}^{(0)};E_{mn})\bigg{)}^{2}-\frac{1}{K^{2}}\bigg{(}\sum_{i=1}^{K}x_{i}(\mathcal{O}^{(0)};E_{mn})\bigg{)}^{2}$
$=(\frac{1}{K}-\frac{1}{K^{2}})\sum_{i=1}^{K}\bigg{(}x_{i}(\mathcal{O}^{(0)};E_{mn})\bigg{)}^{2}-\frac{1}{K^{2}}\sum_{i\neq j}^{K}x_{i}(\mathcal{O}^{(0)};E_{mn})x_{j}(\mathcal{O}^{(0)};E_{mn}).$ (4.9)
[…]
$\sum_{i\neq j}^{K}\langle x_{i}(\mathcal{O}^{(0)};E_{mn})x_{j}(\mathcal{O}^{(0)};E_{mn})\rangle\sim\sum_{i\neq j}^{K}\langle x_{i}(\mathcal{O}^{(0)};E_{mn})\rangle\langle x_{j}(\mathcal{O}^{(0)};E_{mn})\rangle\sim 0.$ (4.10)
This qualitatively can be understood from an analog with random walks and from the nature of the initial distribution (GOE, GUE and the uniform distribution). So after taking the average over initial random operators, we obtain qualitatively
$\langle\sigma_{\mathcal{O}}^{2}\rangle$ $\sim(\frac{1}{K}-\frac{1}{K^{2}})\sum_{i=1}^{K}\bigg{\langle}x_{i}(\mathcal{O}^{(0)};E_{mn})^{2}\bigg{\rangle}.$ (4.11)
If the sample is large enough, we can qualitatively expect $\sigma_{\mathcal{O}}^{2}$ to follow the normal distribution as a result of the central limit theorem.
From the above qualitative analysis, we see that the numbers of random numbers are determined by $\mathcal{N}_{max}$, $K$ and sample size of initial operators. If they are large enough, we would expect the resulting distribution to be normal. In the following computation, we use the choice $\mathcal{N}_{max}=100$ and $K=500$ in consistent with Hashimoto:2023swv , and we average over $5000$ intial operators. This number turns out to be large enough and the resulting distribution function $f_{\sigma^{2}}$ is the normal distribution. (In Appendix C, we check the case $\mathcal{N}_{max}=5$. The resulting distribution $f_{\sigma^{2}}$ is not normal there, and interesting, the phenomenon of separation of $f_{\sigma^{2}}$ for different initial distributions disappears.)
5 Model and Results
For concrete, we study the Sinai billiard which is a typical two-dimensional system of chaos. The Hamiltonian is defined by
$H=p_{x}^{2}+p_{y}^{2}+V(x,y)$ (5.1)
with the potential
$V(x,y)=\begin{cases}0&(x,y)\in\Omega\\ \infty&\mathrm{else}\end{cases}.$ (5.2)
The region $\Omega$ is obtained by cutting out a circle of radius $l$ from an equilateral triangle of side length $L$, as shown in Fig.1. The system is integrable when $l=0$ and chaotic when $0<l<L$. For convenient, the area of $\Omega$ is kept at 1 and the parameter $a\equiv l/L$ is introduced, whose value determines the degree of chaos for the system.
Figure 1: The region $\Omega$ of the potential (5.2) of the billiard. Dirichlet conditions are imposed on the boundaries.
For practical computation in the energy representation, firstly we set the value $\mathcal{N}_{max}$ to truncate the maximum number of energy levels. Here we adopt the convention $\mathcal{N}_{max}=100$ in accordance with Hashimoto:2023swv . Secondly, we pick a random matrix $\mathcal{M}_{\mathcal{O}}$ of size $100\times 100$ from the probability distributions (3.2) and (3.3), which is the matrix of the initial random operator $\mathcal{O}$. Then we compute the Lanczos coefficient $\left\{b_{n}\right\}$ of $\mathcal{K}_{m}(\mathcal{L},\mathcal{M}_{\mathcal{O}})$ by the algorithm in Section 2. To calculate the variance $\sigma^{2}$ of the Lanczos coefficients, we choose $5\mathcal{N}_{max}$ coefficients in consistent with Hashimoto:2023swv . The set of coefficients $\left\{b_{n}\right\}$ is from the algorithm step sizes $500\leq n\leq 1000$, as shown in the Fig.2(a). Note that this selection of the set of $\left\{b_{n}\right\}$ does not qualitatively affect the result of this paper. The effect of selecting the set of $\left\{b_{n}\right\}$ from different algorithm step sizes is discussed in the appendix B. Finally, we repeat this process $5000$ times to get the resulting distribution function $f(\mathcal{O})$.
(a) $b_{n}$ in chaotic case $a=1$.
(b) $b_{n}$ in nonchaotic case $a=0$.
Figure 2: The Lanczos coefficients $\left\{b_{n}\right\}$ for an initial operator selected from GUE. The sets of coefficients between the two red lines ($500\leq n\leq 1000$) are used to calculate the variance $\sigma^{2}$ (2.7).
For the different distributions (GOE, GUE, URE, UIM and UCP) of initial operators, the sampled variances $\sigma^{2}$ are shown in Fig. 3, where $m$ labels the samples. For visual cleanness, we only show $1000$ samples instead of all $5000$ samples. For GOE and GUE, they overlap completely in the nonchaotic case in Fig. 3(b), and separate from each other in the chaotic case in Fig. 3(a). Similary for the uniform distributions, they mix together in the nonchaotic case in Fig. 3(d), and split into two separated groups in the chaotic case in Fig. 3(c), with URE and UIM being one group and UCP being the other.
(a) Sampled variances $\sigma^{2}$ of GOE and GUE in chaotic case $a=1$.
(b) Sampled variances $\sigma^{2}$ of GOE and GUE in nonchaotic case $a=0$.
(c) Sampled variances $\sigma^{2}$ of uniform distributions in chaotic case $a=1$.
(d) Sampled variances $\sigma^{2}$ of uniform distributions in nonchaotic case $a=0$.
Figure 3: Samples of variances $\sigma^{2}$ for GOE, GUE, URE, UIM and UCP. The horizontal axis represents the $m$th sampling. The chaotic $a=1$ and nonchatic case $a=0$ have different behaviors.
The corresponding histogram of the sampled $\sigma^{2}$ is shown in Fig. 4, which shows that the resulting probability distribution resembles the normal distribution. For the Gaussian ensembles, we can see clearly that they are overlapped in the nonchaotic case in Fig. 4(b) and separated in the chaotic case in Fig. 4(a). For the uniform distributions, they are overlapped in the nonchaotic case in Fig. 4(d). In the chaotic case in Fig. 4(c), URE and UIM are still together, but they are separated from UCP.
(a) Histogram of $\sigma^{2}$ of GOE and GUE in the chaotic case $a=1$.
(b) Histogram of $\sigma^{2}$ of GOE and GUE in the nonchaotic case $a=0$.
(c) Histogram of $\sigma^{2}$ of uniform distributions in the chaotic case $a=1$.
(d) Histogram of $\sigma^{2}$ of uniform distributions in the nonchaotic case $a=0$.
Figure 4: The histogram of variances $\sigma^{2}$ after sampling the initial operator 5000 times from GOE, GUE, URE, UIM and UCP. Obviously they resemble normal distributions.
We have shown the Gaussian ensembles and the uniform distributions separately for visual cleanness in the above figures 3 and  4. Their axis values indicate that they can be compared together in one figure. For this purpose, we fit 111 The details of this data fitting is shown in the Appendix A. their histogram to the normal distribution and find the average value $\langle\sigma^{2}\rangle$ and the standard deviation, which characterize the resulting distribution $f_{\sigma^{2}}$. Then the resulting distribution $f_{\sigma^{2}}$ can be shown in a figure where the point is this average value and the error bar is the standard deviation. For all different distributions (GOE, GUE, URE, UIM and UCP) of initial operators, the resulting distributions $f_{\sigma^{2}}$ are shown in Fig. 5, as the system varies from nonchaotic $a=0$ to chaotic $a=1$. We see that when the system is nonchaotic, the data points of all distributions overlap together within the range of error bars. As the system becomes chaotic, different distributions start to separate at $a=0.1$ and the distance of separation is roughly fixed after $a=0.15$. Note that the axis is logarithmic, so the separation is very large. We see that they split into two groups well-separated: one group is of GOE, URE and UIM, the other group is GUE and UCP.
Figure 5: The logarithm plot of the average $\langle\sigma^{2}\rangle$ of resulting normal distributions $f_{\sigma^{2}}$, for all different distributions (GOE, GUE, URE, UIM and UCP) of initial operators. The error bar is the standard deviation. The system dynamics changes from nonchaotic to chaotic, as $a$ varies from 00 to $1$.
6 Discussion
Firstly, we emphasize the importance of the resulting distribution being the normal distribution. We repeat the computation for $\mathcal{N}_{max}=5$ in Appendix C, where the resulting distribution $f_{\sigma^{2}}$ is not normal and there is no separation of $f_{\sigma^{2}}$ in the chaotic case. This suggests that the resulting distributions being normal is essential for their separation behavior in the chaotic case.
Qualitatively we can understand this by the central limit theorem as discussed in Section 4. The resulting distribution being normal means that we have enough random numbers in the sample, so that the statistics is reliable and robust. Luckily, the typical choice of $\mathcal{N}_{max}=100$ in literature is adequate for the normal distribution. It remains an open problem to quantitatively understand the connection between being normal distributions and the separation behavior in the chaotic case.
Secondly, we emphasize that these overlapping-separating behaviors do not depend on the choice of algorithm step-sizes of coefficients $\left\{b_{n}\right\}$. In appendix B, we choose the step-sizes $1000\leq n\leq 1500$ from the Krylov space in Fig. 2. We still get the resulting normal distributions and the overlapping-splitting behaviors. Note that the detailed values of the average and the standard deviation will change, it is the overlapping-separating behavior that does not change. It remains an open problem to quantitatively prove this independence on different choices of algorithm step-sizes.
Thirdly, we discuss the resulting two groups of separation. The GOE, URE and UIM are in one group. We can view GOE, URE and UIM as essentially composed of only real numbers, because the UIM is just real numbers multiplied by the overall factor-the imaginary $i$. The other group is GUE and UCP, which are essentially composed of complex numbers with both real and imaginary parts. So one group is essentially related with one-dimensional random numbers and the other group is related with two-dimensional random numbers. This distinction between real and complex numbers may be related to the nature of the integrability-breaking-term, or to the symmetry of the model. This remains to be one of the open problems.
Finally, for the chaotic case, we compare our separation behavior with a similar behavior-of-separation in RMT Rabinovici:2022beu . For a spin system Rabinovici:2022beu under chaotic dynamics, the Hamiltonian is sampled from RMT ensembles while the initial operator is fixed. For Hamiltonians from GUE and GOE, a separation is obtained in the late-time saturation of the Krylov complexity. Qualitatively, the behavior of late-time saturation of K-complexity are consistent with the behavior of variances $\sigma^{2}$ of the Lanczos coefficients Rabinovici:2021qqt . So it might be possible that the ensemble average of the variance $\sigma^{2}$ over RMT Hamiltonian are also separating in Rabinovici:2022beu , although it is not computed there. If this indeed happen, the separation behavior of the resulting distributions $f_{\sigma^{2}}$ and of the late-time saturation of the Krylov complexity can be viewed as ’dual’ to each other for chaotic systems.
Qualitatively, this ’duality’ can be understood as following: in one side, the Hamiltonian are random and the initial operator is fixed, and in the other side, the Hamiltonian is fixed and the initial operators are random. They are ’equivalent’ when taking the statistics, at least for chaotic systems. From the proof (4.7) of the lemma in Section 4, we see that the Lanczos coefficients and the Krylov basis depend only on the initial operators and energy levels. Both the random energy levels and random initial operators can give random Lanczos coefficients, so intuitively, these statistics can be viewed as dual or equivalent. It remains an open problem to quantitatively prove this connection.
In conclusion, we pick the Sinai billiard and generate random initial operators from given probability distribtions (GOE, GUE, URE, UIM and UCP). Then we find that the resulting distributions of the variance of Lanzcos coefficients are normal distributions. They are overlapped in the nonchaotic case. When the system becomes chaotic, they split into two groups, with one group composed of GOE, URE and UIM, and the other group composed of GUE and UCP. We analyze the reason of being normal and discuss the overlap-separation behavior with open problems.
Acknowledgements.
Wei Fan is supported in part by the National Natural Science Foundation of China under Grant No. 12105121.
Appendix A Check the normal distribution of $\sigma^{2}$ for $\mathcal{N}_{max}=100$
Here we show the data fitting of the histogram in Fig. 4 to the normal distribution. For GOE and GUE, the fitted distribution is shown in Fig. 6. For uniform distributions, the fitted distribution is shown in Fig. 7. The average value and the standard deviation are listed in Table 1, where we compare the values obtained from the sampled data and from the fitted normal distribution.
(a) Nonchaotic case $a=0$ of GOE. Skewness=0.0592927, Kurtosis=3.02995
(b) Nonchaotic case $a=0$ of GUE. Skewness=-0.06833, Kurtosis=3.13681
(c) Chaotic case $a=1$ of GOE. Skewness=0.201339, Kurtosis=3.19744
(d) Chaotic case $a=1$ of GUE. Skewness=0.247936, Kurtosis=2.98557
Figure 6: Fitted normal distribution for GOE and GUE, $\mathcal{N}_{max}=100$.
(a) Chaotic case $a=1$ of UIM. Skewness=0.306957, Kurtosis=3.26486.
(b) Chaotic case $a=1$ of URE. Skewness=0.234259, Kurtosis=2.88038.
(c) Chaotic case $a=1$ of UCP. Skewness=0.207669, Kurtosis=2.97609.
(d) Nonchaotic case $a=0$ of UIM. Skewness=0.0715752, Kurtosis=2.99478.
(e) Nonchaotic case $a=0$ of URE. Skewness=0.0339498, Kurtosis=3.07535.
(f) Nonchaotic case $a=0$ of UCP. Skewness=-0.135062, Kurtosis=3.25632.
x
Figure 7: Fitted normal distribution for uniform distributions, $\mathcal{N}_{max}=100$.
$\mathcal{N}_{max}=100$
$\mu_{0}$ $\sigma_{0}$
nonchaos GOE Fit 0.08205 0.00125033
Data 0.0821079 0.00124627
GUE Fit 0.0818423 0.000799776
Data 0.0818522 0.000818971
URE Fit 0.0819847 0.00122253
Data 0.0819877 0.00122892
UIM Fit 0.0822362 0.001221
Data 0.0822584 0.0012246
UCP Fit 0.0818089 0.000796413
Data 0.0817893 0.000813977
chaos GOE Fit 0.000360084 0.0000311866
Data 0.000361536 0.0000312431
GUE Fit 0.000239301 0.0000196699
Data 0.00024081 0.0000194523
URE Fit 0.000313361 0.0000277979
Data 0.000315455 0.0000278845
UIM Fit 0.000314084 0.0000270829
Data 0.000315694 0.0000267367
UCP Fit 0.000210981 0.0000167099
Data 0.000211866 0.0000168488
Table 1: Average $\mu_{0}$ and standard deviation $\sigma_{0}$ for the data fitting, $\mathcal{N}_{max}=100$. ’Data’ means the values computed from the sampled data. ’Fit’ means the values of the fitted normal distribution.
Appendix B Different choice of algorithm step sizes for $b_{n}$
For the overlapping-separating behavior, here we show its independence on the sets of coefficients used in computing $\sigma^{2}$. Now we choose the Lanczos coefficients $\left\{b_{n}\right\}$ from step-sizes $1000\leq n\leq 1500$.
Firstly, we compute the variance $\sigma^{2}$ (2.7) for GOE and GUE, where the first $1000$ samples are shown in Fig. 8. We see that GOE and GUE mix together in the nonchaotic case and separate from each other in the chaotic case.
(a) Sampled variances $\sigma^{2}$ of GOE and GUE in chaotic case $a=1$.
(b) Sampled variances $\sigma^{2}$ of GOE and GUE in nonchaotic case $a=0$.
Figure 8: The first $1000$ samples of variances $\sigma^{2}$ for GOE and GUE. Here the Lanczos coefficients $\left\{b_{n}\right\}$ from step-sizes $1000\leq n\leq 1500$ are used.
Then, the histogram of all $5000$ samples are shown in Fig. 9. Obviously they still resemble normal distributions and are overlapped in the nonchaotic case and well-separated in the chaotic case.
(a) Histogram of $\sigma^{2}$ of GOE and GUE in the chaotic case $a=1$.
(b) Histogram of $\sigma^{2}$ of GOE and GUE in the chaotic case $a=1$.
Figure 9: The histogram of variances $\sigma^{2}$ after sampling the initial operator 5000 times from GOE and GUE. The Lanczos coefficients $\left\{b_{n}\right\}$ from step-sizes $1000\leq n\leq 1500$ are used. They are overlapped in the nonchaotic case and well-separated in the chaotic case.
Finally, the fitted normal distributions are shown in Fig. 10. So we still get the resulting normal distributions and the overlapping-splitting behaviors, although the detailed values of the average and the standard deviation change.
(a) Nonchaotic case $a=0$ of GOE. Skewness=-0.113535, Kurtosis=2.94238.
(b) Nonchaotic case $a=0$ of GUE. Skewness=-0.156533, Kurtosis=3.16475.
(c) Chaotic case $a=1$ of GOE. Skewness=0.107514, Kurtosis=3.01354.
(d) Chaotic case $a=1$ of GUE. Skewness=0.0881284, Kurtosis=3.06162.
x
Figure 10: Fitted normal distribution for GOE and GUE, $\mathcal{N}_{max}=100$. The Lanczos coefficients $\left\{b_{n}\right\}$ from step-sizes $1000\leq n\leq 1500$ are used.
Appendix C Check the normal distribution of $\sigma^{2}$ for $\mathcal{N}_{max}=5$
Here we study the case of $\mathcal{N}_{max}=5$. Then the sets of coefficients $b_{n}$ are chosen from step-sizes $151<n<$ 15. We repeat the computation for GOE and GUE, and show that there is neither normal distributions nor separating behaviors.
The histogram of all $5000$ samples are shown in Fig. 11. Obviously they are not normal distributions and are not separated in the chaotic case.
(a) Histogram of $\sigma^{2}$ of GOE and GUE in the chaotic case $a=1$.
(b) Histogram of $\sigma^{2}$ of GOE and GUE in the nonchaotic case $a=0$.
Figure 11: The histogram of variances $\sigma^{2}$ after sampling the initial operator 5000 times from GOE and GUE, $\mathcal{N}_{max}=5$. Obviously they are not normal distributions.
Anyway, if we force to fit the histogram to normal distributions, the results are shown Fig. 12. From the ridiculous value of skewness and kurtosis, we understand that the histogram can not be the normal distribution. Also, from the values listed in Table 2, we see that there is large discrepancy between values from the sample data and values from the fitted distribution. So for $\mathcal{N}_{max}=5$, the resulting distributions are not normal and there is no separating behaviors between GOE and GUE.
(a) $N=5$ chaos, Gaussian U ensembles. Skewness=1.19843, Kurtosis=5.18562.
(b) $N=5$ nonchaos, Gaussian U ensembles. Skewness=4.30801, Kurtosis=50.7359.
(c) $N=5$ chaos, Gaussian O ensembles. Skewness=1.6231, Kurtosis=7.05429.
(d) $N=5$ nonchaos, Gaussian O ensembles. Skewness=5.83862, Kurtosis=73.6201.
Figure 12: Data fitting for GOE and GUE, $\mathcal{N}_{max}=5$. The large value of skewness and kurtosis means that they can not be normal.
$\mathcal{N}_{max}=5$
$\mu_{0}$ $\sigma_{0}$
nonchaos GOE Fit 0.128068 0.099812
Data 0.188669 0.178039
GUE Fit 0.0967919 0.0465464
Data 0.119184 0.0752851
URE Fit 0.119756 0.0814372
Data 0.169944 0.146747
UIM Fit 0.206881 0.220034
Data 0.467013 0.602757
UCP Fit 0.0678911 0.0397609
Data 0.0877107 0.0616032
chaos GOE Fit 0.127619 0.0872543
Data 0.170123 0.112041
GUE Fit 0.104014 0.0521167
Data 0.121707 0.0610372
URE Fit 0.112995 0.0772997
Data 0.150244 0.0991289
UIM Fit 0.20724 0.0973048
Data 0.253492 0.14551
UCP Fit 0.0688293 0.0407356
Data 0.0853616 0.0509862
Table 2: Average $\mu_{0}$ and standard deviation $\sigma_{0}$ for the data fitting, $\mathcal{N}_{max}=5$. There is large discrepancy between values computed from the sampled data and values of the fitted normal distribution.